OpenStudy (anonymous):

find the exact value of -8cos(11 pi/6)

OpenStudy (anonymous):

\[\Huge \color{Red}{\Pi=180^° } \]

OpenStudy (anonymous):

\[\huge \color{red}{\text{if } \pi=180,\frac{ \pi }{ 6 }}=?\]

OpenStudy (anonymous):